# Local analysis of iterative reconstruction from discrete generalized Radon transform data in the plane

### v1

@@ 3.3. GRT data and the reconstruction algorithm
(3.21) $\begin{split}&\Upsilon_{l}(r)=\int_{\mathbb{R}}h_{l}(u)\int_{u_{l}}^{u_{l}+r}R(t)\text{d}t\text{d}u,\ \vec{\Theta}_{0}:=\frac{H_{x}^{\prime}(x_{0})}{|H_{x}^{\prime}(x_{0})|},\\ &h_{l}(u)=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(x_{0},\alpha_{l})+u)\text{d}s,\ u_{l}=\frac{u}{|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|},\\ &R(t)=\mathcal{F}_{\lambda\to t}^{-1}\bigg{(}\bigg{[}1+\frac{\kappa}{2\pi\textstyle\sum_{l}\rho_{l}}|\lambda|^{3}\bigg{]}^{-1}\bigg{)},\ \rho_{l}:=\frac{W^{2}(x_{0},y_{l})|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|}{|\Delta_{\Phi}(x_{0},\alpha_{l})|}.\end{split}$
@@ 4. Behavior of $\mathcal{R}f$ near its singular support
(4.1) $\Gamma=\text{sing\hskip 1.0ptsupp}(\mathcal{R}f),\ WF(\mathcal{R}f)\subset N^{*}\Gamma=\mathcal{C}\circ N^{*}\mathcal{S}.$
Here $N^{*}\mathcal{S}$ is the conormal bundle of $\mathcal{S}$, and similarly for $\Gamma$ [52, section I.6]. In particular, $\mathcal{R}f$ is smooth away from $\Gamma$. In this section we describe the leading singular behavior of $\mathcal{R}f$ near $\Gamma$.
Using a partition of unity and the linearity of $\mathcal{R}$, we can assume without loss of generality that $\mathcal{S}_{y}$, $y\in\Gamma$, is tangent to $\mathcal{S}$ only at one point.
Fix any $\tilde{y}\in\Gamma$, and let $\tilde{x}\in\mathcal{S}$ be the corresponding point of tangency. Let $V$ be a sufficiently small neighborhood of $\tilde{y}$. Substitute (3.7) into (3.1). The asymptotics as $\lambda\to\infty$ of the resulting integral with respect to $x$ is computed by the stationary phase method [53, Chapter VIII, eqs. (2.14)–(2.20)]
(4.2) $\begin{split}(\mathcal{R}f)(y)=&\frac{1}{2\pi}\int_{\mathbb{R}}\tilde{G}(y,\lambda)\text{d}\lambda,\ y\in V,\\ \tilde{G}(y,\lambda):=&\frac{1}{2\pi}\int_{\mathbb{R}}\int_{\mathcal{U}}\tilde{f}(x,\lambda)e^{i\lambda H(x)}W(x,y)e^{i\nu(p-\Phi(x,\alpha))}\text{d}x\text{d}\nu\\ =&\frac{|\lambda|}{2\pi}\int_{\mathbb{R}}\int_{\mathcal{U}}W(x,y)\tilde{f}(x,\lambda)e^{i\lambda[\nu_{1}(p-\Phi(x,\alpha))+H(x)]}\text{d}x\text{d}\nu_{1}\\ =&\left(\tilde{f}(x_{*},\lambda)\frac{W(x_{*},y)}{|\det M(y)|^{1/2}}\left(\frac{2\pi}{|\lambda|}\right)^{1/2}e(-\text{sgn}\lambda/2)+\tilde{R}(y,\lambda)\right)\\ &\quad\times e^{i\lambda H(x_{*})},\ |\lambda|\geq 1,\ \tilde{R}\in S^{-3/2}(\mathcal{V}\times\mathbb{R}),\ x_{*}=x_{*}(y),\end{split}$
for some $\tilde{R}$ and $0<\delta\ll 1$. Here $e(t):=e(i\pi t/2)$, and
(4.3) $M(y)=\begin{pmatrix}0&(-\Phi_{x}^{\prime}(x,\alpha))^{T}\\ -\Phi_{x}^{\prime}(x,\alpha)&H_{xx}^{\prime\prime}(x,\alpha)-\nu_{1}\Phi_{xx}^{\prime\prime}(x,\alpha)\end{pmatrix},\ \nu_{1}=\nu_{1*}(y),x=x_{*}(y),$
@@
By construction, $x_{*}(\tilde{y})=\tilde{x}$. As before, the existence and local uniqueness of the solution follows from assumptions 3.2(1,3).
The following result is proven in appendix B.1.
@@
Suppose assumptions 3.13.3 are satisfied. Let $\mathcal{S}_{y}$ be tangent to $\mathcal{S}$ at $x\in\mathcal{S}$. Let $\varkappa_{\mathcal{S}}(x)$ and $\varkappa_{\mathcal{S}_{y}}(x)$ denote the curvatures of $\mathcal{S}$ and $\mathcal{S}_{y}$ at $x$, respectively (see (3.11)). One has
(4.5) $\begin{split}\det M(y)=&(\varkappa_{\mathcal{S}}(x)-\varkappa_{\mathcal{S}_{y}}(x)){|\Phi_{x}^{\prime}(x,\alpha)|^{2}}{|H_{x}^{\prime}(x)|}>0,\\ \text{sgn}M(y)=&-1.\end{split}$
@@
(4.7) $\begin{split}\psi(\alpha,P(\alpha))=&\partial_{p}H(x_{*}(\alpha,p))|_{p=P(\alpha)}=\frac{|H_{x}^{\prime}(x_{*}(y))|}{|\Phi_{x}^{\prime}(x_{*}(y),\alpha)|}>0,\\ (\alpha,P(\alpha))\in&V.\end{split}$
@@
It is clear that instead of finding $P(\alpha)$ through the intermediary function $x_{*}(y)$, we can find $P(\alpha)$ by solving
(4.8) $H(x(\alpha))=0,\ H_{x}^{\prime}(x(\alpha))=\nu_{1}(\alpha)\Phi_{x}^{\prime}(x(\alpha),\alpha),\ (\alpha,P(\alpha))\in V,$
for $\nu_{1}(\alpha)$, $x(\alpha)$ and setting $P(\alpha):=\Phi(x(\alpha),\alpha)$. Then $\nu_{1}(\alpha)\equiv\nu_{1*}(\alpha,P(\alpha))$. Since $x(\alpha)\in\mathcal{S}$, we can say that $x(\alpha)$ and $P(\alpha)$ are associated with a segment of $\mathcal{S}$ (see Figure 1). The size of the segment depends on the domain of $x(\alpha)$. If we have in mind one specific $\tilde{x}\in\mathcal{S}$, then we informally say that $x(\alpha)$ and $P(\alpha)$ are associated with $\tilde{x}$ (i.e., with a segment of $\mathcal{S}$ containing $\tilde{x}$).
@@
By (4.2)–(4.6), $\mathcal{R}f$ can be written as
(4.9) $(\mathcal{R}f)(y)=\frac{1}{2\pi}\int_{\mathbb{R}}\tilde{\upsilon}(y,\lambda)e^{i\lambda(p-P(\alpha))}\text{d}\lambda,\ y\in V,$
@@
(4.10) $\begin{split}&\tilde{\upsilon}(y,\lambda)=\tilde{\upsilon}_{0}(y)e(-3/2)(\lambda-i0)^{-3/2}+\tilde{R}(y,\lambda),\ |\lambda|\geq 1;\\ &\tilde{\upsilon}\in S^{-3/2}(V\times\mathbb{R});\ \tilde{R}\in S^{-5/2}(V\times\mathbb{R});\\ &\tilde{\upsilon}_{0}(y)=(2\pi)^{1/2}\frac{W(x_{*}(y),y)\tilde{f}_{0}(x_{*}(y))\psi(y)^{1/2}}{|\det M(y)|^{1/2}}\in C^{\infty}(V).\end{split}$
@@
where $\tilde{G}$ is defined in (4.2). To obtain the formula for $\tilde{\upsilon}_{0}$ we use that the leading order term in $\tilde{\upsilon}$ is homogeneous of degree $-3/2$ and the distribution $(\lambda-i0)^{a}$ [14, equations (28) and (30), p. 336].
@@
(4.12) sing $\text{sing\hskip 1.0ptsupp}(\mathcal{R}f)\cap V\subset\Gamma:=\{(\alpha,p)\in V:p=P(\alpha)\}.$
All of the above properties of $\mathcal{R}f$ follow from the FIO calculus [53, Section VIII.5], but for our purposes it is more convenient to provide their explicit derivation (4.2)–(4.11).
By appealing to a partition of unity and the linearity of the GRT, the following global result is proven in appendix B.2.
@@
Suppose assumptions 3.13.4 are satisfied. Let $g=\mathcal{R}f$. Then $g,g_{\epsilon}\in C_{0}^{1/2}(\mathcal{V})$ and $\|g_{\epsilon}\|_{C^{1/2}(\mathcal{V})}<c$, $0<\epsilon\ll 1$. Also,
(4.13) $\begin{split}&g(y)=\textstyle\sum_{l}a_{l}(y)(p-P_{l}(\alpha))_{+}^{1/2}+\Delta g(y),\\ &a_{l}(y)=4\pi^{1/2}\tilde{\upsilon}_{0,l}(y),\ a_{l}\in C_{0}^{\infty}(\mathcal{V}),\ \Delta g\in C_{0}^{3/2}(\mathcal{V}),\end{split}$
and $\|\Delta g_{\epsilon}\|_{C^{3/2}(\mathcal{V})}<c$ for all $0<\epsilon\ll 1$. Here, for each $l$,
  1. (1)
    $\text{supp}\,a_{l}$ is contained in a small neighborhood of some $y_{l}\in\Gamma$;
  2. (2)
    $p=P_{l}(\alpha)$ is a local solution to $H(x_{*}(\alpha,p))=0$ for $p$ in terms of $\alpha$, which is defined for all $\alpha$ such that $(\alpha,P_{l}(\alpha))\in\text{supp}\,a_{l}$; and
  3. (3)
    $\tilde{\upsilon}_{0,l}(y)$ is computed as in (4.10).

### v2

@@ 3.3. GRT data and the reconstruction algorithm
(3.22) $\begin{split}&\Upsilon_{l}(r)=\int_{\mathbb{R}}h_{l}(u)\int_{u_{l}}^{u_{l}+r}R(t)\text{d}t\text{d}u,\ \vec{\Theta}_{0}:=\frac{\text{d}_{x}H(x_{0})}{|\text{d}_{x}H(x_{0})|},\\ &h_{l}(u)=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(x_{0},\alpha_{l})+u)\text{d}s,\ u_{l}=\frac{u}{|\text{d}_{x}\Phi(x_{0},\alpha_{l})|},\\ &R(t)=\mathcal{F}_{\lambda\to t}^{-1}\bigg{(}\bigg{[}1+\frac{\kappa}{2\pi\textstyle\sum_{l}\nu_{l}}|\lambda|^{3}\bigg{]}^{-1}\bigg{)},\ \nu_{l}:=\frac{W^{2}(x_{0},y_{l})|\text{d}_{x}\Phi(x_{0},\alpha_{l})|}{|\Delta_{\Phi}(x_{0},\alpha_{l})|}.\end{split}$
Taken together, Theorems 3.6 and 3.7 provide an easy to use characterization of the reconstructed image in an $O(\epsilon)$-size neighborhood of an edge. This characterization is in terms of universal functions $\Upsilon_{0}$ and $\Upsilon_{l}$, which are independent of $f$ and can be easily precomputed. The only contribution from $f$ is $\Delta f(x_{0})$, the value of the jump of $f$ at $x_{0}$, which enters the formulas (3.18) and (3.21) as a multiplicative factor. Although the proofs are quite extensive, the final result is satisfying due to its simple form and straightforward applicability.
3.4. Example
To briefly illustrate the above theorems, we consider the classical Radon transform in the plane that integrates over the lines $\mathcal{S}_{(\alpha,p)}=\{x\in\mathbb{R}^{2}:\vec{\alpha}\cdot x=p\}$, $\vec{\alpha}:=(\cos\alpha,\sin\alpha)$. Then $\mathcal{I}_{\alpha}=[0,2\pi)\sim S^{1}$, $\mathcal{I}_{p}=\mathbb{R}$, $\Phi(x,\alpha)=\vec{\alpha}\cdot x$, $W(x,y)\equiv 1$, and
(3.23) $\begin{split}&\text{d}_{x}\Phi(x,\alpha)=\vec{\alpha},\ |\text{d}_{x}\Phi(x,\alpha)|=1,\ \Phi_{\alpha}^{\prime}(x,\alpha)=\vec{\alpha}^{\perp}\cdot x,\\ &\Delta_{\Phi}=\det\begin{pmatrix}\vec{\alpha}\\ \vec{\alpha}^{\perp}\end{pmatrix}=1,\ \vec{\alpha}^{\perp}:=(-\sin\alpha,\cos\alpha).\end{split}$
Pick any $x_{0}\in\mathcal{S}$. Let $\vec{\alpha}_{0}$ be perpendicular to $\mathcal{S}$ at $x_{0}$. Strictly speaking, there are two data points $y_{1}=(\alpha_{0},\vec{\alpha}_{0}\cdot x_{0})$ and $y_{2}=(\alpha_{0}+\pi,-\vec{\alpha}_{0}\cdot x_{0})$ such that $\mathcal{S}_{y_{l}}$ is tangent to $\mathcal{S}$ at $x_{0}$. Hence we should use Theorem 3.7.
From (3.22) and (3.23), $u_{l}=u$, $\nu_{l}=1$, and
(3.24) $\begin{split}&\Upsilon_{l}(r)=\int_{\mathbb{R}}h_{l}(u)\int_{u}^{u+r}R(t)\text{d}t\text{d}u,\\ &h_{l}(u)=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\vec{\alpha}_{l}^{\perp}\cdot x_{0}+u)\text{d}s,\ l=1,2,\\ &R(t)=\mathcal{F}_{\lambda\to t}^{-1}\bigg{(}\big{[}1+\frac{\kappa/2}{2\pi}|\lambda|^{3}\big{]}^{-1}\bigg{)}.\end{split}$
By Assumption 3.4(2), $\varphi_{\alpha}$ is even and $\vec{\alpha}_{1}^{\perp}\cdot x_{0}=-\vec{\alpha}_{2}^{\perp}\cdot x_{0}$. Therefore, $h_{1}(u)\equiv h_{2}(u)$ and (cf. (3.21))
(3.25) $\frac{\textstyle\sum_{l}\nu_{l}\Upsilon_{l}(r)}{\textstyle\sum_{l}\nu_{l}}=\int_{\mathbb{R}}h_{1}(u)\int_{u}^{u+r}R(t)\text{d}t\text{d}u.$
This is the formula we would have obtained by formally using Theorem 3.6 and replacing $\kappa$ with $\kappa/2$. The result is not surprising, as $\mathcal{S}_{y_{1}}$ and $\mathcal{S}_{y_{2}}$ define the same tangent line.
See Section 10 for an example involving the GRT that integrates over circles. The functions $\Upsilon_{l}$ have been used to compute the predicted behavior of the reconstruction (the DTB function), which is shown as a blue curve in Figure 3.
@@ 4. Behavior of $\mathcal{R}f$ near its singular support
(4.1) $\Gamma=\text{sing\hskip 1.0ptsupp}\hat{f},\ \text{WF}(\hat{f})\subset N^{*}\Gamma=\mathcal{C}\circ N^{*}\mathcal{S}.$
Here $N^{*}\mathcal{S}$ is the conormal bundle of $\mathcal{S}$, and similarly for $\Gamma$ [55, section I.6]. In particular, $\hat{f}$ is smooth away from $\Gamma$. In this section we describe the leading singular behavior of $\hat{f}$ near $\Gamma$.
Fix any $\tilde{y}\in\Gamma$, and let $\tilde{x}\in\mathcal{S}$ be the corresponding point of tangency, see Figure 1. Let $V$ be a sufficiently small neighborhood of $\tilde{y}$. Using a partition of unity, the linearity of $\mathcal{R}$, and Assumption 3.3(2), we can assume without loss of generality that $\mathcal{S}_{y}$, $y\in\Gamma\cap V$, is tangent to $\mathcal{S}$ only at one point. We underscore that the main goal of this section is to study the singularities of $\mathcal{R}f$. The way these singularities are mapped into the reconstructed image is studied in Sections 7, 8, and 9.
Substitute (3.7) into (3.1). The asymptotics as $\lambda\to\infty$ of the resulting integral with respect to $x$ is computed by the stationary phase method [56, Chapter VIII, eqs. (2.14)–(2.20)] using the properties of the Hessian at the stationary point established in Lemma 4.1 below:
(4.2) $\begin{split}\hat{f}(y)=&\frac{1}{2\pi}\int_{\mathbb{R}}\tilde{G}(y,\lambda)\text{d}\lambda,\ y\in V,\\ \tilde{G}(y,\lambda):=&\frac{1}{2\pi}\int_{\mathbb{R}}\int_{\mathcal{U}}\tilde{f}(x,\lambda)e^{i\lambda H(x)}W(x,y)e^{i\nu(p-\Phi(x,\alpha))}\text{d}x\text{d}\nu\\ =&\frac{|\lambda|}{2\pi}\int_{\mathbb{R}}\int_{\mathcal{U}}W(x,y)\tilde{f}(x,\lambda)e^{i\lambda[\nu_{1}(p-\Phi(x,\alpha))+H(x)]}\text{d}x\text{d}\nu_{1}\\ =&\left(\tilde{f}(x_{*},\lambda)\frac{W(x_{*},y)}{|\det M(y)|^{1/2}}\left(\frac{2\pi}{|\lambda|}\right)^{1/2}e(-\text{sgn}\lambda/2)+\tilde{R}(y,\lambda)\right)\\ &\quad\times e^{i\lambda H(x_{*})},\ |\lambda|\geq 1,\ \tilde{R}\in S^{-3/2}(\mathcal{V}\times\mathbb{R}),\ x_{*}=x_{*}(y),\end{split}$
for some $\tilde{R}$. Here $e(t):=\exp(i\pi t/2)$, and
(4.3) $\begin{split}&M(y)=\begin{pmatrix}0&-\text{d}_{x}\Phi(x,\alpha)\\ (-\text{d}_{x}\Phi(x,\alpha))^{T}&\text{d}_{x}^{2}\big{[}H(x,\alpha)-\nu_{1}\Phi(x,\alpha)\big{]}\end{pmatrix},\\ &\nu_{1}=\nu_{1*}(y),\ x=x_{*}(y),\end{split}$
@@
By construction, $x_{*}(\tilde{y})=\tilde{x}$. As before, the existence, local uniqueness, and smoothness of the solution follows from Assumptions 3.2(1,3) and 3.3(2).
Given a nondegenerate symmetric matrix $M$, let $n_{+}$ and $n_{-}$ denote the number of positive and negative eigenvalues of $M$, respectively. Then the signature of $M$, denoted $\text{sgn}\,M$, is the difference $n_{+}-n_{-}$. The following result is proven in Appendix B.1.
@@
Suppose Assumptions 3.13.3 are satisfied. Let $\mathcal{S}_{y}$ be tangent to $\mathcal{S}$ at $x$. Let $\varkappa_{\mathcal{S}}(x)$ and $\varkappa_{\mathcal{S}_{y}}(x)$ denote the curvatures of $\mathcal{S}$ and $\mathcal{S}_{y}$ at $x$, respectively (see (3.11)). One has
(4.5) $\begin{split}\det M(y)=&(\varkappa_{\mathcal{S}}(x)-\varkappa_{\mathcal{S}_{y}}(x)){|\text{d}_{x}\Phi(x,\alpha)|^{2}}{|\text{d}_{x}H(x)|}>0,\\ \text{sgn}\,M(y)=&-1.\end{split}$
@@
(4.7) $\begin{split}\psi(\alpha,P(\alpha))=&\partial_{p}H(x_{*}(\alpha,p))|_{p=P(\alpha)}=\frac{|\text{d}_{x}H(x_{*}(y))|}{|\text{d}_{x}\Phi(x_{*}(y),\alpha)|}>0,\\ (\alpha,P(\alpha))\in&V.\end{split}$
@@
Instead of finding $P(\alpha)$ via the intermediary function $x_{*}(y)$, we can find $P(\alpha)$ by locally solving
(4.8) $H(x(\alpha))=0,\ \text{d}_{x}H(x(\alpha))=\nu_{1}(\alpha)\text{d}_{x}\Phi(x(\alpha),\alpha),$
for $\nu_{1}(\alpha)$, $x(\alpha)$ and setting $P(\alpha):=\Phi(x(\alpha),\alpha)$. Then $\nu_{1}(\alpha)\equiv\nu_{1*}(\alpha,P(\alpha))$. Since $x(\alpha)\in\mathcal{S}$, we can say that $P(\alpha)$ and the segment of $\Gamma$ it defines are associated with a segment of $\mathcal{S}$ (see Figure 1). If we have in mind one specific $\tilde{x}\in\mathcal{S}$, then we informally say that $P(\alpha)$ and the corresponding segment of $\Gamma$ are associated with $\tilde{x}$ (i.e., with a segment of $\mathcal{S}$ containing $\tilde{x}$).
@@
By (4.2)–(4.6), $\hat{f}$ can be written as
(4.9) $\hat{f}(y)=\frac{1}{2\pi}\int_{\mathbb{R}}\tilde{\upsilon}(y,\lambda)e^{i\lambda(p-P(\alpha))}\text{d}\lambda,\ y\in V,$
@@
(4.10) $\begin{split}&\tilde{\upsilon}(y,\lambda)=\tilde{\upsilon}_{0}(y)e(-3/2)(\lambda-i0)^{-3/2}+\tilde{R}(y,\lambda),\ |\lambda|\geq 1;\\ &\tilde{\upsilon}\in S^{-3/2}(V\times\mathbb{R});\ \tilde{R}\in S^{-5/2}(V\times\mathbb{R});\\ &\tilde{\upsilon}_{0}(y)=(2\pi)^{1/2}\frac{W(x_{*}(y),y)\Delta f(x_{*}(y))\psi(y)^{1/2}}{|\det M(y)|^{1/2}}\in C^{\infty}(V).\end{split}$
@@
where $\tilde{G}$ is defined in (4.2). To obtain the formula for $\tilde{\upsilon}_{0}$ we use that the leading order term in $\tilde{\upsilon}$ is homogeneous of degree $-3/2$ and the definition of the distribution $(\lambda-i0)^{a}$ [17, equations (28) and (30), p. 336].
@@
(4.12) sing $\text{sing\hskip 1.0ptsupp}\hat{f}\cap V\subset\Gamma:=\{(\alpha,p)\in V:p=P(\alpha)\}.$
All of the above properties of $\hat{f}$ follow from the FIO calculus [56, Section VIII.5], but for our purposes it is more convenient to provide their explicit derivation (4.2)–(4.11).
As mentioned above, Assumptions 3.1(1,3) and 3.3(1) imply $\text{supp}\,\hat{f}$ is compact. Consider an open cover of $\text{supp}\,\hat{f}$ by sufficiently small open sets $V_{l}:=I_{\alpha}^{l}\times I_{p}^{l}$, $1\leq l\leq L$, for some large $L$. Here $I_{\alpha}^{l}$ and $I_{p}^{l}$ are intervals. Using a partition of unity subordinate to this cover, we can consider one such set, $V_{l}$. Generally, the intersection $\Gamma\cap V_{l}$ is a single smooth curve segment. If $\Gamma$ self-intersects in $V_{l}$, we formally duplicate $V_{l}$ as many times as needed to make sure each intersection $\Gamma\cap V_{l}$ is identified with no more than one smooth curve segment, denoted $\Gamma_{l}$.
Let $p=P_{l}(\alpha)$, $\alpha\in I_{\alpha}^{l}$, be the local parameterization for $\Gamma_{l}$. Unless stated otherwise, we consider only one set from the partition, denoted $V=I_{\alpha}\times I_{p}$. The corresponding $P_{l}(\alpha)$ is denoted $P(\alpha)$. Also, we denote $J_{\alpha}:=\{j_{1}\in\mathbb{Z}:\alpha_{j_{1}}\in I_{\alpha}\}$, and similarly for $J_{p}$. There are $O(1/\epsilon)$ points in $J_{\alpha}$ and $J_{p}$.
The following result is proven in Appendix B.2.
@@
Suppose Assumptions 3.13.4 are satisfied. Pick any $\chi_{V}\in C_{0}^{\infty}(V)$. Let $g:=\chi_{V}\hat{f}$. Then $g,g_{\epsilon}\in C_{0}^{1/2}(V)$ and $\|g_{\epsilon}\|_{C^{1/2}(V)}<c$, $0<\epsilon\ll 1$. Also,
(4.13) $\begin{split}&g(y)=a_{0}(y)(p-P(\alpha))_{+}^{1/2}+\Delta g(y),\\ &a(y)=2\pi^{-1/2}\tilde{\upsilon}_{0}(y),\ a\in C_{0}^{\infty}(V),\ \Delta g\in C_{0}^{3/2}(\mathcal{V}),\end{split}$
and $\|\Delta g_{\epsilon}\|_{C^{3/2}(V)}<c$ for all $0<\epsilon\ll 1$, where $\tilde{\upsilon}_{0}(y)$ is given in (4.10). Here
(4.14) $g_{\epsilon}(y):=\textstyle\sum_{j\in V}\varphi_{\epsilon}(y-y_{j})g(y_{j}),$
and $\Delta g_{\epsilon}$ is defined analogously, see (3.14).
If $\Gamma\cap V=\varnothing$, we assume in (4.13) that $a_{0}\equiv 0$.